# fftvis: A Non-Uniform Fast Fourier Transform Based Interferometric Visibility SimulatorA

### v1

@@ 2.2 Non-Uniform Fast-Fourier Transform Evaluation of RIME
where $X_{i}=\max_{j}|x^{(i)}_{j}|$ represents the maximum extent of the source coordinates in dimension $i$, and $S_{i}=\max_{k}|s^{(i)}_{k}|$ represents the maximum extent of the target baseline coordinates in units of wavelengths in that dimension. The parameter $\sigma>1$ is a grid oversampling factor (typically $\sigma=2$) used to mitigate aliasing effects. The product of these terms $X_{i}S_{i}$ captures the core requirement of the NUFFT: a larger extent in either the spatial domain ($X_{i}$) or the Fourier domain ($S_{i}$) requires a larger number of grid points ($n_{i}$) to maintain accuracy.
It is important to note that the overall intermediate grid size, $N_{\rm grid}=\prod_{i}^{d}n_{i}$ (defined more formally below), is thus driven by these maximum extents, $X_{i}$ and $S_{i}$, rather than by the sheer number of sources ($N_{\rm sources}$) or baselines ($N_{\rm bls}$) contained within those extents. Effectively, the grid must be fine enough to resolve the highest Fourier variations ($S_{i}$) across the full spatial extent ($X_{i}$) of the input, plus an additional margin determined by the kernel-width ($w$) and oversampling factor ($\sigma$) to achieve the target precision ($\epsilon$). While larger $X_{i}S_{i}$ products increase the grid size and thus the cost of the FFT, finufft mitigates this by automatically translating both $\{\mathbf{x}_{j}\}$ and $\{\mathbf{s}_{k}\}$ to be centered around zero, minimizing $X_{i}$ and $S_{i}$ and hence reducing the grid size. However, when simulating wide-field visibilities (large $X_{i}$) or including baselines with large $|\mathbf{b}|/\lambda$ (large $S_{i}$), the required grid size can still become substantial. Conversely, if either the angular extent of the sky being simulated is small (small $X_{i}$), or the array’s extent in wavelengths is limited (small $S_{i}$), $N_{\rm grid}$ can remain more manageable, which is beneficial for NUFFT performance regardless of $N_{\rm sources}$ or $N_{\rm bls}$ within those compact extents. In such cases, the performance may depend sensitively on the geometry of the problem.
@@
This computational scaling reflects the three-stage evaluation of the RIME, in which $N_{\rm sources}$ point sources are convolved onto the intermediate grid with a kernel of size $w^{d}$, then Fourier transformed via an FFT, which has a computational scaling of $\mathcal{O}\left(N_{\rm grid}\log N_{\rm grid}\right)$, and finally interpolated to $N_{\rm bls}$ baselines with the Fourier transform of the gridding kernel, which also has size $w^{d}$. This provides a substantial speedup over a direct $O(N_{\rm sources}N_{\rm bls})$ evaluation for simulation cases where the number of baselines or the number of sources in the sky model is large. The NUFFT advantage over direct summation is particularly maximized when the intermediate FFT grid size ($N_{\rm grid}$) remains manageable. This occurs not only when the physical array size in units of wavelengths is small (small $S_{i}$), but also when simulating a compact region of the sky (small $X_{i}$), as $N_{\rm grid}$ depends on these extents rather than directly on $N_{\rm sources}$ or $N_{\rm bls}$. However, most high redshift 21 cm applications require a full-sky, horizon-to-horizon sky model. In this case, we can write the scaling of the grid size in more familiar terms,
@@ 2.2.1 Evaluating the Measurement Equation for Gridded Arrays
The Type 1 NUFFT can also be adapted for arrays where their antenna positions form a regular but non-Cartesian lattice (e.g. a hexagonal grid). In these instances, a linear transformation $L$ is chosen to map the physical baselines $\mathbf{b}_{\rm phys}$ into a set of gridded baselines, $\mathbf{b}_{\rm gridded}=L\mathbf{b}_{\rm phys}$, which are, by design, arranged on a regular Cartesian grid. To ensure that the phase term ($-2\pi i\nu\mathbf{b}_{\rm phys}\cdot\mathbf{x}_{j}/c$) in the visibility sum is accurately computed, the original source coordinates $\mathbf{x}_{j}$ must also be transformed to $\mathbf{x}_{j}^{\prime}=(L^{T})^{-1}\mathbf{x}_{j}$. The Type 1 NUFFT then calculates the visibilities values corresponding to these gridded baseline configurations $\mathbf{b}_{\rm gridded}$ using the transformed source coordinates $\mathbf{x}_{j}^{\prime}$ as input.
@@
This can offer a significant performance advantage over the general Type 3 NUFFT’s $\mathcal{O}(N_{\rm sources}w^{d}+N_{\rm grid}\log N_{\rm grid}+N_{\rm bls}w^{d})$ scaling, particularly if $N_{\rm unif}$ is substantially smaller than $N_{\rm grid}$ (the Type 3 intermediate grid size from Equation 5). This occurs when the array is regular, but very sparse. $N_{\rm grid}$’s size is strongly influenced by the product of the source extent $X_{i}$ and the maximum extent of the target modes $S_{i}$ (where $S_{i}$ relates to the maximum baseline components scaled by $\nu/c$). In contrast, $N_{\rm unif}$ reflects the number of points needed to represent the specific set of gridded baseline configurations of the array. If this count is smaller than $N_{\rm grid}$ (especially if $X_{i}$ is large, which tends to inflate $N_{\rm grid}$ for Type 3), the Type 1 NUFFT provides a more efficient route for simulating visibilities.
@@ 2.2.2 Considerations for Non-Flat Arrays
The NUFFT framework and its performance characteristics, including the determination of $N_{\rm grid}$ (Equation 5) and the overall scaling involving $w^{d}$, apply generally to $d$ dimensions. For arrays that are effectively co-planar, or where baseline $w$-components (the component parallel to the pointing direction) are negligible, simulations can often utilize a 2-dimensional NUFFT ($d=2$) which is typically less computationally intensive. However, when an array contains antenna with significant non-coplanar positions, its baseline vectors $\mathbf{b}/\lambda=(u,v,w)$ will include substantial $w$-components. To accurately model the source-baseline geometry in these scenarios, the dot product $\nu\mathbf{b}\cdot\mathbf{x}_{j}/c$ in the visibility equation must be treated in its full 3-dimensional form.
Using a 3D NUFFT (that is, setting $d=3$ in the Type 3 transform algorithm as used by fftvis) has significant consequences for the computational cost, impacting each stage of the NUFFT process detailed earlier. Firstly, the gridding of $N_{\rm sources}$ onto the intermediate grid and the final interpolation from this grid to $N_{\rm bls}$ baselines now involve a 3-dimensional convolution kernel. This means that their operational costs, which scale with $w^{d}$, will now scale with $w^{3}$ rather than $w^{2}$. Secondly, and often more importantly, the size of the intermediate FFT grid, $N_{\rm grid}$, becomes a 3D product, $N_{\rm grid}=n_{x}n_{y}n_{z}$. Each $n_{i}$ (for $i\in\{x,y,z\}$) is determined by Equation 5, so $n_{z}$ will depend on the maximum extent of the product of the baseline $w$-components (contributing to $S_{z}$) and the corresponding source coordinate ($X_{z}$, related to the $n_{j}-1$ term in direction cosines). Even if $n_{z}$ (the grid dimension along the $w$-axis) is smaller than $n_{x}$ or $n_{y}$, its inclusion substantially increases the total number of points in $N_{\rm grid}$ compared to a 2D grid of $n_{x}n_{y}$. As a result, the FFT stage, with complexity $\mathcal{O}(N_{\rm grid}\log N_{\rm grid})$, becomes much more computationally intensive due to this significantly larger $N_{\rm grid}$.
@@ 2.2.3 Assumption of Identical Antenna Beams
where $N_{\rm beam}$ is the number of unique beam patterns present within the array. If we assume that the NUFFT step of the simulation dominates the total runtime, we expect this will increase the simulation time by a factor of $N_{\rm NUFFT}$. In contrast, the matrix-based approach used by matvis has a more favorable scaling with $N_{\rm beams}$ since the core visibility computation depends on the number of antennas rather than the number of unique beam patterns. To maintain computational efficiency, the current implementation of fftvis assumes the beam model supplied by the user is shared by all antennas. While this constraint remains well-justified for many practical simulation scenarios, certain validation and analysis tasks require support for per-antenna beams (Orosz et al., 2019; Wilensky et al., 2024; Kim et al., 2022, 2023). Future releases of fftvis will include support for heterogeneous, antenna-dependent beam configurations.
@@ 2.3 Coordinate Rotation
By default, fftvis uses the hybrid method where full astrometric corrections via the astropy library are applied only periodically (at user-specified intervals) to pre-correct source coordinates. Between these corrections, computationally efficient rigid body rotations, dependent on local sidereal time and latitude, transform these pre-corrected equatorial coordinates to the horizontal (ENU) frame at each time step. In Kittiwisit et al. 2025, the authors demonstrate that this hybrid transformation approach reproduces the results of full-precision coordinate evaluations with excellent accuracy for snapshot observations.
However, because the transformation matrix $\mathbf{R}(t)$ is constructed assuming a fixed celestial sphere, errors due to neglected time-dependent astrometric effects (e.g., precession, proper motion, and nutation) increase with time since the reference epoch. This results in periodic errors in the simulated visibilities on a 24-hour sidereal cycle relative to the high-precision simulations produced by pyuvsim, which utilizes astropy for its coordinate rotations and transformations. From a simulation standpoint, the primary advantage of this scheme lies in its computational efficiency: once equatorial positions are pre-corrected, subsequent evaluations require only matrix multiplications that are easily vectorized and parallelized. However, this same simplification limits its long-term astrometric accuracy, particularly for simulations requiring precision modeling for calibration purposes. For many 21 cm cosmology use cases, especially those involving short observations, the hybrid method achieves a practical balance between computational cost and astrometric accuracy.
@@ 2.4 Beam Interpolation
Another component in achieving accurate simulations is the precise representation of antenna beam patterns evaluated at continuously evolving sky coordinates. This requirement presents particular challenges as sources drift across the field of view, necessitating beam evaluation at arbitrary sky coordinates where empirical beam measurements or simulations may not have explicitly been covered. Our beam interpolation strategy follows the well-established techniques used by the matvis codebase backed by pyuvdata, ensuring that our approach takes advantage of thoroughly validated computational methods for beam handling and interpolation. The framework accommodates two principal classes of antenna beam models:
@@
For beam data management and interpolation operations, we integrate the AnalyticBeam and UVBeam classes from the pyuvdata package (Hazelton et al., 2017), maintained by the Radio Astronomy Software Group (RASG) [^2] https://radioastronomysoftwaregroup.github.io/. This framework supports various beam representation formats, including regularly gridded beam models, HEALPix-formatted beam distributions, and CST-simulated electromagnetic beam patterns. The UVBeam class provides multiple interpolation functions with configurable order parameters, including bilinear, bicubic, and quintic interpolation schemes operating on altitude-azimuth coordinate systems. These methods use the computational interfaces scipy.interpolate.RectBivariateSpline and scipy.ndimage.map_coordinates to achieve high-precision beam evaluation at arbitrary source positions.
There are a few practical considerations, emphasized in Kittiwisit et al. 2025, that become especially important in high-accuracy simulations. One concerns the choice of interpolation scheme: when modeling visibilities for calibration or foreground subtraction, where sub-percent level deviations can become important, higher-order interpolation methods are recommended to maintain the accuracy of the beam response. However, the authors did note that for validation purposes, bilinear interpolation did not introduce appreciable systematic effects in the 21 cm power spectrum estimate, indicating that lower-order interpolation schemes are sufficient for applications that do not demand full physical realism. Another relates to the coordinate system used for beam definition. Many E-field beam models exhibit phase discontinuities at the zenith in at least one polarization component. To handle these correctly, the beam should be defined on a spherical grid centered at the zenith, allowing interpolation routines to trace the field structure smoothly and avoid introducing artifacts associated with non-polar or Cartesian grids.

### v2

@@ 2.2 Non-Uniform Fast-Fourier Transform Evaluation of RIME
[…]
It is important to note that the overall intermediate grid size, $N_{\rm grid}=\prod_{i}^{d}n_{i}$ (defined more formally below), is thus driven by these maximum extents, $X_{i}$ and $S_{i}$, rather than by the sheer number of sources ($N_{\rm sources}$) or baselines ($N_{\rm bls}$) contained within those extents. Effectively, the grid must be fine enough to resolve the highest Fourier variations ($X_{i}$) across the full spatial extent ($S_{i}$) of the input, plus an additional margin determined by the kernel-width ($w$) and oversampling factor ($\sigma$) to achieve the target precision ($\epsilon$). While larger $X_{i}S_{i}$ products increase the grid size and thus the cost of the FFT, finufft mitigates this by automatically translating both $\{\mathbf{x}_{j}\}$ and $\{\mathbf{s}_{k}\}$ to be centered around zero, minimizing $X_{i}$ and $S_{i}$ and hence reducing the grid size. However, when simulating wide-field visibilities (large $S_{i}$) or including baselines with large $|\mathbf{b}|/\lambda$ (large $X_{i}$), the required grid size can still become substantial. Conversely, if either the angular extent of the sky being simulated is small (small $S_{i}$), or the array’s extent in wavelengths is limited (small $X_{i}$), $N_{\rm grid}$ can remain more manageable, which is beneficial for NUFFT performance regardless of $N_{\rm sources}$ or $N_{\rm bls}$ within those compact extents. In such cases, the performance may depend sensitively on the geometry of the problem.
@@
This computational scaling reflects the three-stage evaluation of the RIME, in which $N_{\rm sources}$ point sources are convolved onto the intermediate grid with a kernel of size $w^{d}$, then Fourier transformed via an FFT, which has a computational scaling of $\mathcal{O}\left(N_{\rm grid}\log N_{\rm grid}\right)$, and finally interpolated to $N_{\rm bls}$ baselines with the Fourier transform of the gridding kernel, which also has size $w^{d}$. This provides a substantial speedup over a direct $O(N_{\rm sources}N_{\rm bls})$ evaluation for simulation cases where the number of baselines or the number of sources in the sky model is large. The NUFFT advantage over direct summation is particularly maximized when the intermediate FFT grid size ($N_{\rm grid}$) remains manageable. This occurs not only when the physical array size in units of wavelengths is small (small $X_{i}$), but also when simulating a compact region of the sky (small $S_{i}$), as $N_{\rm grid}$ depends on these extents rather than directly on $N_{\rm sources}$ or $N_{\rm bls}$. However, most high redshift 21 cm applications require a full-sky, horizon-to-horizon sky model. In this case, we can write the scaling of the grid size in more familiar terms,
@@ 2.2.1 Evaluating the Measurement Equation for Gridded Arrays
The Type 1 NUFFT can also be adapted for arrays where their antenna positions form a regular but non-Cartesian lattice (e.g. a hexagonal grid). In these instances, a linear transformation $\mathbf{L}$ is chosen to map the physical baselines $\mathbf{b}_{\rm phys}$ into a set of gridded baselines, $\mathbf{b}_{\rm gridded}=\mathbf{L}\mathbf{b}_{\rm phys}$, which are, by design, arranged on a regular Cartesian grid. To ensure that the phase term ($2\pi i\nu\mathbf{b}_{\rm phys}\cdot\mathbf{s}_{j}/c$) in the visibility sum is accurately computed, the original source coordinates $\mathbf{s}_{j}$ must also be transformed to $\mathbf{s}_{j}^{\prime}=(\mathbf{L}^{T})^{-1}\mathbf{s}_{j}$. The Type 1 NUFFT then calculates the visibilities values corresponding to these gridded baseline configurations $\mathbf{b}_{\rm gridded}$ using the transformed source coordinates $\mathbf{s}_{j}^{\prime}$ as input.
@@
This can offer a significant performance advantage over the general Type 3 NUFFT’s $\mathcal{O}(N_{\rm sources}w^{d}+N_{\rm grid}\log N_{\rm grid}+N_{\rm bls}w^{d})$ scaling, particularly if $N_{\rm unif}$ is substantially smaller than $N_{\rm grid}$ (the Type 3 intermediate grid size from Equation 5). This occurs when the array is regular, but very sparse. $N_{\rm grid}$’s size is strongly influenced by the product of the source extent $S_{i}$ and the maximum extent of the target modes $X_{i}$ (where $X_{i}$ relates to the maximum baseline components scaled by $\nu/c$). In contrast, $N_{\rm unif}$ reflects the number of points needed to represent the specific set of gridded baseline configurations of the array. If this count is smaller than $N_{\rm grid}$ (especially if $S_{i}$ is large, which tends to inflate $N_{\rm grid}$ for Type 3), the Type 1 NUFFT provides a more efficient route for simulating visibilities.
By default, the fftvis codebase automatically inspects the user-supplied antenna positions to determine whether they lie on a lattice, and therefore whether a computationally cheaper Type 1 NUFFT can be substituted for the general Type 3 case. This procedure computes the set of unique baseline separations in each Cartesian direction and evaluates whether these separations can be expressed as integer multiples of a common fundamental spacing within a user-defined numerical tolerance. If this condition is satisfied, fftvis identifies the corresponding lattice vectors, constructs the linear transformation matrix $\mathbf{L}$ that maps the physical baseline vectors $\mathbf{b}_{\rm phys}$ onto a regular grid ($\mathbf{b}_{\rm gridded}=\mathbf{L}\,\mathbf{b}_{\rm phys}$), and applies the appropriate inverse transpose $(\mathbf{L}^{T})^{-1}$ to the input sky coordinates. The transformation is fully automated, and no requires no manual pre-processing of antenna or source coordinates by the user. When the array geometry does not meet the lattice criteria, either because the spacings are irregular beyond the specified tolerance or the configuration is inherently non-gridded, fftvis defaults to the Type 3 NUFFT. This ensures numerical correctness for arbitrary array layouts while exploiting the Type 1 speedup whenever possible.
@@ 2.2.2 Considerations for Non-Flat Arrays
The NUFFT framework and its performance characteristics, including the determination of $N_{\rm grid}$ (Equation 5) and the overall scaling involving $w^{d}$, apply generally to $d$ dimensions. For arrays that are effectively co-planar, or where baseline $w$-components (the component parallel to the pointing direction) are negligible, simulations can often utilize a 2-dimensional NUFFT ($d=2$) which is typically less computationally intensive. However, when an array contains antenna with significant non-coplanar positions, its baseline vectors $\mathbf{b}/\lambda=(u,v,w)$ will include substantial $w$-components. To accurately model the source-baseline geometry in these scenarios, the dot product $\nu\mathbf{b}\cdot\mathbf{s}_{j}/c$ in the visibility equation must be treated in its full 3-dimensional form.
Using a 3D NUFFT (that is, setting $d=3$ in the Type 3 transform algorithm as used by fftvis) has significant consequences for the computational cost, impacting each stage of the NUFFT process detailed earlier. Firstly, the gridding of $N_{\rm sources}$ onto the intermediate grid and the final interpolation from this grid to $N_{\rm bls}$ baselines now involve a 3-dimensional convolution kernel. This means that their operational costs, which scale with $w^{d}$, will now scale with $w^{3}$ rather than $w^{2}$. Secondly, and often more importantly, the size of the intermediate FFT grid, $N_{\rm grid}$, becomes a 3D product, $N_{\rm grid}=n_{x}n_{y}n_{z}$. Each $n_{i}$ (for $i\in\{x,y,z\}$) is determined by Equation 5, so $n_{z}$ will depend on the maximum extent of the product of the baseline $w$-components (contributing to $X_{z}$) and the corresponding source coordinate ($S_{z}$, related to the $n_{j}-1$ term in direction cosines). Even if $n_{z}$ (the grid dimension along the $w$-axis) is smaller than $n_{x}$ or $n_{y}$, its inclusion substantially increases the total number of points in $N_{\rm grid}$ compared to a 2D grid of $n_{x}n_{y}$. As a result, the FFT stage, with complexity $\mathcal{O}(N_{\rm grid}\log N_{\rm grid})$, becomes much more computationally intensive due to this significantly larger $N_{\rm grid}$.
@@ 2.2.3 Assumption of Identical Antenna Beams
where $N_{\rm beam}$ is the number of unique beam patterns present within the array. If we assume that the NUFFT step of the simulation dominates the total runtime, we expect this will increase the simulation time by a factor of $N_{\rm NUFFT}$. In contrast, the matrix-based approach used by matvis has a more favorable scaling with $N_{\rm beams}$ since the core visibility computation depends on the number of antennas rather than the number of unique beam patterns. To maintain computational efficiency, the current implementation of fftvis assumes the beam model supplied by the user is shared by all antennas. While this constraint remains well-justified for many practical simulation scenarios, certain validation and analysis tasks require support for per-antenna beams (2019MNRAS.487..537O; 2024RASTI...3..400W; 2022ApJ...941..207K; 2023ApJ...953..136K). Future releases of fftvis will include support for heterogeneous, antenna-dependent beam configurations.
@@ 2.3 Coordinate Rotation
By default, fftvis uses the hybrid method where full astrometric corrections via the astropy library are applied only periodically (at user-specified intervals) to pre-correct source coordinates. Between these corrections, computationally efficient rigid body rotations, dependent on local sidereal time and latitude, transform these pre-corrected equatorial coordinates to the horizontal (ENU) frame at each time step. In 2025RASTI...4....1K, the authors demonstrate that this hybrid transformation approach reproduces the results of full-precision coordinate evaluations with excellent accuracy for snapshot observations.
However, because the transformation matrix $\mathbf{R}(t)$ is constructed assuming a fixed celestial sphere, errors due to neglected time-dependent astrometric effects (e.g., precession, proper motion, and nutation) increase with time since the reference epoch. Specifically, slow-varying effects like precession and nutation are applied only once during the astropy-based pre-correction at the reference time, not during the per-timestep rigid-body rotations. Because these astrometric effects change very slowly, their error contribution is negligible for observations taken at or near this reference time. These simplifications result in periodic errors in the simulated visibilities on a 24-hour sidereal cycle relative to the high-precision simulations produced by pyuvsim, which utilizes astropy for its coordinate rotations and transformations. The error is quantified in 2025RASTI...4....1K, where the authors show that for snapshot observations, the hybrid method matches high-precision simulations with fractional residuals of $10^{-10}$.
From a simulation standpoint, the primary advantage of this scheme lies in its computational efficiency: once equatorial positions are pre-corrected, subsequent evaluations require only matrix multiplications that are easily vectorized and parallelized. However, this same simplification limits its long-term astrometric accuracy, particularly for simulations requiring precision modeling for calibration purposes. For many 21 cm cosmology use cases, especially those involving short observations, the hybrid method achieves a practical balance between computational cost and astrometric accuracy. For longer simulations requiring higher precision, 2025RASTI...4....1K recommend mitigating this error by splitting the observation into shorter chunks and refreshing the full astrometric correction at a regular interval. In their analysis, the authors found that updating these astrometric corrections every $\sim 15$ minutes of simulation time kept fractional visibility errors to less than $\sim 10^{-5}$. As this interval is user-specified in fftvis, the $\sim 15$ minute recommendation from the matvis paper provides guidance on how to determine the necessary update interval to balance computational cost and the required astrometric accuracy.
@@ 2.4 Beam Interpolation
Another component in achieving accurate simulations is the precise representation of antenna beam patterns evaluated at continuously evolving sky coordinates. This requirement presents particular challenges as sources drift across the field of view, necessitating beam evaluation at arbitrary sky coordinates where empirical beam measurements or simulations may not have explicitly been covered. Our beam interpolation strategy follows the well-established techniques used by the matvis codebase backed by pyuvdata (2017JOSS....2..140H; 2025JOSS...10.7482K), ensuring that our approach takes advantage of thoroughly validated computational methods for beam handling and interpolation. The framework accommodates two principal classes of antenna beam models:
@@
For beam data management and interpolation operations, we integrate the AnalyticBeam and UVBeam classes from the pyuvdata package, maintained by the Radio Astronomy Software Group (RASG) [^3] https://radioastronomysoftwaregroup.github.io/. This framework supports various beam representation formats, including regularly gridded beam models, HEALPix-formatted beam distributions, and CST-simulated electromagnetic beam patterns. The UVBeam class provides multiple interpolation functions with configurable order parameters, including bilinear, bicubic, and quintic interpolation schemes operating on altitude-azimuth coordinate systems. These methods use the computational interfaces scipy.interpolate.RectBivariateSpline and scipy.ndimage.map_coordinates to achieve high-precision beam evaluation at arbitrary source positions.
There are a few practical considerations, emphasized in 2025RASTI...4....1K, that become especially important in high-accuracy simulations. One concerns the choice of interpolation scheme: when modeling visibilities for calibration or foreground subtraction, where sub-percent level deviations can become important, higher-order interpolation methods are recommended to maintain the accuracy of the beam response. However, the authors did note that for validation purposes, bilinear interpolation did not introduce appreciable systematic effects in the 21 cm power spectrum estimate, indicating that lower-order interpolation schemes are sufficient for applications that do not demand full physical realism. Another relates to the coordinate system used for beam definition. Many E-field beam models exhibit phase discontinuities at the zenith in at least one polarization component. To handle these correctly, the beam should be defined on a spherical grid centered at the zenith, allowing interpolation routines to trace the field structure smoothly and avoid introducing artifacts associated with non-polar or Cartesian grids.